# GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v3

@@ Appendix E Experiment Details
E.1. Further introduction of chosen baselines
@@
  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019). Entailment then carry out classification based on these entailment pairs and convert the text classification task into a binary classification task, for detecting whether the entailment pair formed is correct. When all entailment pairs of a particular text are considered to be incorrect, Entailment will directly give up the classification of this text. In this paper, we consider this classification as a hallucination.
@@
  • NaiveRAG (Gao et al., 2023): NaiveRAG acts as a foundational baseline of current RAG models. When indexing, it stores text segments of the labeled texts in a vector database using text embeddings. When querying, NaiveRAG generates query texts’ vectorized representations to retrieve side information based on the highest similarity in their embeddings.
  • Propositionizer (Chen et al., 2023): Propositionizer applies a fine-tuned LLM to convert side information into atomic expressions, namely propositions, to facilitate more fine-grained information retrieval than NaiveRAG.
Compression RAG Models
  • LongLLMLingua (Jiang et al., 2024a): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
  • GraphRAG (Edge et al., 2024): GraphRAG is a graph-Based RAG model that employs the LLM to extract texts’ entities and relations, which are then represented as nodes and edges in the weighted graph. GraphRAG then aggregates nodes into communities, and generates a comprehensive community report to encapsulate global information from texts.
  • LightRAG (Guo et al., 2024): LightRAG skips the GraphRAG’s formulation of graph communities and directly retrieve nodes or pathes from created graphs. It calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping from keywords to graph nodes and retrieve all triples involved these nodes.
@@
0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed weighted graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.
0:  $\hat{\mathcal{Y}}^{r}_{t}$: The candidate type retrieved for text $t$.
@@
2:  Let $\mathcal{V}_{term}=\mathcal{V}_{exist}^{t}\cap\mathcal{V}(MST)$ be the terminal nodes in $MST$
@@
5:     Find the shortest path $P_{uv}$ in $MST$ from $u$ to $v$
@@
E.2. Hyperparameter and Hardware Settings
In each round of our experiments, we handle NN and RAG-based models differently for training and indexing. For NN-based models, we train them using all labeled data from the current round and all previous rounds. For RAG-based models, we index the labeled data of the current round to the information source from the previous round. The RAG-based model is initialized from scratch only in the first round.
After training or indexing in each round, we test the models using the testing data from the current round and all previous rounds. For GORAG, which employs an online indexing mechanism, we first test its performance with online indexing on the current round’s testing data. We then test its performance on the testing data from all previous rounds without online indexing. This approach allows us to study the effect of online indexing on the performance of later round data when applied to previous round data.
For the hyperparameter settings, we train the Entailment model with the RoBERTa-large PLM for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-6}$, and the Adam optimizer (Kingma, 2014), following the exact setting in Entailment’s original paper. For LongLLMLingua, we test the compression rate within ${0.75,0.8,0.85}$ and select 0.8, as it achieves the best overall classification accuracy. For GraphRAG and LightRAG, we use their local search mode, as it achieves the highest classification accuracy on both the WOS and Reuters datasets. For GORAG and all RAG-based baselines, we use LLaMA3 (Touvron et al., 2023) as the LLM backbone unless otherwise specified.
@@ E.2. Hyperparameter and Hardware Settings
E.3. Further Analysis based on Precision and F1
According to Table 8 and Table 2, although Entailment achieves the second best accuracy than GORAG on WOS dataset, it has a high hallucination rate. This is because the ground truth label of Entailment’s entailment pairs are greatly unbalanced, hence make the model easy to overfit, and tend to predict all entailment pairs of a particular text as incorrect under few-shot settings.
(a) GORAG’s candidate selection algorithm returns candidate labels adaptively based on keywords extracted from the tesing texts.
(b) GORAG’s candidate selection algorithm only returns the ground truth label of the query text.
Figure 6. Two case of GORAG successfully generate candidate labels to help text classification on WOS dataset.
E.4. GORAG with different backbone LLMs
Table 7. Evaluation on different LLM backbones.
Dataset Model Round
R1 R2 R3 R4
WOS GORAG LLaMA3 0.4862 0.4649 0.4814 0.4210
GORAG Qwen2.5 0.5101 0.4839 0.4823 0.4235
In Table 7, we present the results of GORAG with different backbone LLMs, we test GORAG with more advanced Qwen2.5-7B (Qwen Team, 2024), the result demonstates the GORAG’s performance can be further improved with more advanced LLM backbones.
E.5. Case Study
In this section, to illustrate the strength of GORAG’s adaptive retrieval and candidate label generation procedure, we dig into two testing cases select from WOS dataset where GORAG’s adaptive retrieval helps to reduce the LLM input length and benefit text classification. As shown in Figure 6(a), for the case whose ground truth label is Algorithm Design, the candidate label retrieved by GORAG are Computer programming and Algorithm design. As a result, GORAG’s adaptive retrieval algorithm successfully filter the target label set to only contains two candidate labels, and successfully cut down the target label number from at most over 100 to only 2 candidate labels that it consided as the possible labels for the text, significantlly reduce the LLM’s input length and mitigate the lost-in-the-middle issue.
On the other hand, for the case in Figure 6(b) whose ground truth label is Electrical Circuits, since the ground truth label name Electrical Circuits already exists in the text and being extracted as keyword, GORAG’s adaptive retrieval algorithm is almost certain to classify this text into class Electrical Circuits, hence GORAG’s adaptive retrieval and candidate type generation algorithm would only select the only candidate label retrieved from the constructed graph and input to the LLM.
Table 8. The macro recall score and hallucination rate (Hall. Rate) on WOS dataset.
Metrics Category Model Round
R1 R2 R3 R4
Recall NN-based Entailment 0.3502 0.3475 0.3908 0.3021 0.3415 0.3442 0.2327 0.2406 0.2432 0.1789 0.1792 0.1837
Long Context RAG NaiveRAG 0.3324 0.3417 0.3398 0.1615 0.1652 0.1650 0.0943 0.1013 0.0872 0.0568 0.0805 0.0628
Propositionizer 0.1032 0.1131 0.1114 0.0947 0.1210 0.1187 0.1352 0.1298 0.1254 0.1373 0.1345 0.1402
Compression RAG LongLLMLingua 0.3139 0.3105 0.3258 0.1349 0.1388 0.1350 0.0862 0.0837 0.0872 0.0616 0.0610 0.0628
GraphRAG 0.3791 0.3646 0.3732 0.1535 0.1519 0.1557 0.1009 0.1014 0.1015 0.0902 0.0940 0.0912
LightRAG 0.3462 0.3441 0.3519 0.1544 0.1507 0.1466 0.1241 0.1257 0.1226 0.1129 0.1180 0.1101
GORAG 0.4387 0.4659 0.4874 0.3074 0.3472 0.3483 0.2371 0.2513 0.2445 0.1960 0.1970 0.2013
Hal. Rate NN-based Entailment 54.53% 58.96% 51.63% 44.76% 36.62% 37.76% 32.57% 28.14% 27.35% 29.68% 26.86% 24.13%
Long Context RAG NaiveRAG 3.00% 2.32% 2.63% 6.45% 6.16% 6.28% 6.93% 7.46% 7.52% 9.96% 9.50% 9.83%
Propositionizer 6.85% 5.88% 6.97% 7.40% 6.52% 6.09% 5.97% 6.44% 6.30% 5.78% 6.11% 5.52%
Compression-RAG LongLLMLingua 6.50% 6.29% 5.63% 16.45% 16.05% 16.28% 19.63% 20.47% 20.52% 21.96% 21.35% 21.83%
GraphRAG 2.40% 2.23% 2.07% 6.03% 5.97% 6.18% 7.20% 7.58% 7.03% 6.90% 6.39% 6.49%
LightRAG 1.70% 1.24% 1.37% 2.59% 2.78% 3.21% 4.66% 4.55% 4.58% 5.89% 5.80% 5.73%
GORAG 0.66% 1.16% 0.74% 0.30% 0.68% 0.65% 0.41% 0.57% 0.68% 0.40% 0.43% 0.64%
Table 9. The macro recall score and hallucination rate (Hall. Rate) on IFS-Rel dataset.
Metrics Category Model Round
R1 R2 R3 R4
Recall NN-based Entailment 0.3191 0.6015 0.5515 0.2172 0.3043 0.3330 0.1225 0.1802 0.2081 0.1093 0.1267 0.1490
Long Context RAG NaiveRAG 0.4039 0.4194 0.4292 0.2320 0.2009 0.2115 0.1147 0.1163 0.1097 0.0871 0.0895 0.0912
Propositionizer 0.1053 0.1146 0.1104 0.1124 0.1201 0.1197 0.0824 0.1029 0.0902 0.0744 0.0717 0.0799
Compression RAG LongLLMLingua 0.4426 0.4625 0.4044 0.1879 0.1559 0.2158 0.0947 0.0906 0.1113 0.0698 0.0822 0.0696
GraphRAG 0.4264 0.4139 0.4208 0.2116 0.2008 0.2047 0.1213 0.1179 0.1185 0.0906 0.0857 0.0883
LightRAG 0.4221 0.3985 0.4382 0.2344 0.2150 0.2429 0.1270 0.1313 0.1276 0.0902 0.0926 0.0918
GORAG 0.4588 0.4681 0.5000 0.2348 0.2175 0.2405 0.1279 0.1295 0.1290 0.0950 0.0886 0.0919
Hal. Rate NN-based Entailment 8.10% 15.63% 4.69% 26.41% 19.14% 6.25% 26.56% 17.86% 6.67% 6.48% 5.90% 3.28%
Long Context RAG NaiveRAG 1.56% 4.25% 1.78% 3.05% 2.97% 3.13% 3.02% 3.70% 3.54% 4.02% 4.18% 3.63%
Propositionizer 7.82% 7.34% 11.25% 5.78% 6.40% 6.72% 4.84% 6.09% 7.03% 8.13% 8.59% 6.72%
Compression-RAG LongLLMLingua 3.59% 2.97% 6.56% 2.12% 6.23% 5.47% 6.36% 6.00% 3.50% 3.29% 5.65% 8.88%
GraphRAG 4.25% 1.56% 1.25% 2.66% 3.98% 2.89% 3.39% 3.75% 3.80% 3.91% 3.24% 3.09%
LightRAG 3.28% 2.66% 2.03% 2.73% 2.97% 2.27% 1.98% 1.56% 1.67% 2.46% 2.23% 2.30%
GORAG 0.31% 0.63% 0.16% 1.78% 2.81% 2.58% 1.30% 0.78% 1.09% 1.48% 1.02% 1.25%

### v4

@@ Appendix E Experiment Details
E.1. Introduction of chosen baselines
As shown in Table 2, we compare GORAG’s performance with 7 baselines from 3 technical categories for few-shot experiments, they are NN-based Entailment (Xia et al., 2021), Long Context-based NaiveRAG (Gao et al., 2023), LongLLMLingua (Jiang et al., 2024), and Graph-based GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2025), and HippoRAG2 (gutiérrez2025ragmemory).
@@ E.1. Introduction of chosen baselines
  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019).
@@
  • NaiveRAG (Gao et al., 2023): It acts as a foundational baseline of current RAG models. When indexing, it stores text segments in a vector database using text embeddings. When querying, retrieves the side information based on their vector similarity to the query.
  • LongLLMLingua (Jiang et al., 2024): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
Graph-based RAG Models
  • GraphRAG (Edge et al., 2024): It employs LLM for graph construction. When retrieval, it aggregates nodes into communities w.r.t. the query, and generates community reports to encapsulate global information from texts.
  • LightRAG (Guo et al., 2025): It calculates the vector similarity between query extracted entities and graph nodes, achieving a one-to-one mapping from entities to graph nodes, then retrieve these nodes’ graph neighborhood.
  • HippoRAG2 (gutiérrez2025ragmemory): It employs LLM for graph construction, and also add the text chunks as nodes in the graph. As a result, it directly retrieves Top-k text chunks based on their respective nodes’ PPR score.
@@
[…]
@@
2: Let $\mathcal{V}_{term}=\mathcal{V}_{exist}^{t}\cap\mathcal{V}(MST)$ be the terminal nodes in MSTMST
@@
5:  Find the shortest path $P_{uv}$ in MSTMST from $u$ to $v$
@@
E.2. Hardware Settings
@@
E.3. Case Study
In this section, we dig into a test case from the CAD social media abuse detection dataset to demonstrate the ability of GORAG to provide graph-grounded clues, to explain its classification result.
As shown in Figure 5, the keyword entity feminists and sexual predators are extracted from the input unlabeled text and mapped to the graph. GORAG then generates a Minimum Cost Steiner Tree through these mapped entities, which also spans through the candidate label Insult. The structure and involved nodes of this generated Steiner Tree (edges in red) serve as an explicit clue, helping the LLM to classify the text correctly as Insulting Content.
Figure 5. A case from the CAD dataset, where the generated Steiner Tree serves as an explicit clue, helping the LLM to classify the text correctly.